# What are they talking about? Benchmarking Large Language Models for Knowledge-Grounded Discussion Summarization

### v1

@@ 7 Conclusion
In this study, we introduce the KGDS task, which aims to create observer-centric summaries by integrating shared background knowledge with discussions. We establish the first benchmark for KGDS and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. Our evaluation of leading LLMs reveals significant challenges in background summary retrieval, generation, and opinion summary integration, with even the most advanced models achieving less than average performance across both patterns. These findings highlight the need for future advancements in LLM capabilities for coarse-grained retrieval, fine-grained generation and knowledge integration.
@@
Our study presents several limitations that merit discussion and future exploration:
Domain Generalization.
Knowledge-grounded discussions and their confusing summaries are prevalent in both open-domain and private scenarios. Our benchmark is centered on open-domain news discussions, which may limit its applicability to private scenarios (e.g., internal meetings, medical consultations). However, shared background knowledge and the related discussions in private contexts are challenging to obtain and are less common and representative than our open benchmark. Future work will focus on validating the performance of LLMs in handling KGDS across diverse private scenarios.
Language Diversity.
@@
Our KGDS benchmark is developed with careful consideration of ethical implications. The news articles used as shared background knowledge are publicly accessible through Google News and have been carefully reviewed to ensure they do not contain sensitive or harmful information. Explicit consent was obtained from all expert participants involved in discussion construction and data annotation, with a clear explanation of the research purpose and data usage protocols. Our evaluation methodology prioritizes factual accuracy and opinion fidelity to minimize the risk of hallucination propagation in real-world applications. Potential misuse risks, such as generating misleading summaries through improper background-discussion combinations, are mitigated through technical safeguards in our released code and explicit usage guidelines. Researchers using this benchmark should adhere to responsible AI principles, especially when applying similar techniques to sensitive domains like healthcare or legal discussions. We provide the annotation and evaluation costs as:
Annotation Cost for KGDS BenchMark.
Our annotation process is multi-step, fine-grained, and highly complex. For each sample in our benchmark, the average annotation time is hours (including news reading and understanding, discussion construction, back-supporting paragraphs, key back-supporting atomic facts, back-nonsupporting atomic facts, and clear atomic opinions annotation). We pay each annotation expert a wage of $ per hour (above the minimum wage standard), and the total annotation cost for the KGDS benchmark is $ ( hours per sample $ per hour experts per sample samples $).
Evaluation Cost for KGDS BenchMark.
The evaluation cost of API calls is detailed in Table 3.
@@ Acknowledgement
This research was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081) and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18). We also thank the anonymous reviewers for their constructive feedback that helped improve this work.
Appendix A Overall Performance Visualization
Figures 6 and 8 show the overall performance of EBS-AOS and ABS-AOS. Figures 7 and 9 illustrate the performance gap and average performance.
Appendix B KGDS BenchMark Statistics
@@
Among the paragraphs from original news articles, the expert annotation consistency rate is . A total of paragraphs are retained for the final news articles, while are identified as boundary paragraphs and removed. Of the retained, are annotated as back-supporting and as back-nonsupporting.
@@
Among the atomic fact units decomposed from the back-supporting paragraphs, are consistently annotated as key back-supporting atomic facts, as non-key back-supporting atomic facts, and as intra-repetitive atomic facts.
Among the atomic fact units from the back-nonsupporting paragraphs, are automatically annotated as back-nonsupporting atomic facts, as masked conflicting atomic facts, and as intra-repetitive atomic facts.
@@
A total of clear atomic opinions are expert manually annotated from discussions and their corresponding new articles. Of these, contain clarified implicit references, while do not require clarification. Within the opinions, there are a total of written clarified implicit references.
Appendix C Detailed EBS Evaluation Metrics
Let , , and denote the sets of paragraph indices for LLM-extracted EBS, back-supporting, and back-nonsupporting paragraphs, respectively. The automatic paragraph index matching function is:
where and represent whether the paragraph index is present in .
The BSP Recall, Precision, and F1-score are defined as follows:
Appendix D Prompts and Instructions
D.1 Unified Parameter Settings
In this work, All evaluated LLMs use consistent parameter settings for prompts and instructions: max_token= and temperature=. No other default parameters are modified.
D.2 Atomic Fact Decomposition Instruction
Refer to Figure 11.
D.3 Structured Prompts
D.4 Self-Reflection Instructions
D.5 Feedback Judgment Instruction
Refer to Figure 16.
D.6 Verification Prompts
D.7 Error Detection Instruction
Refer to Figure 18.
@@ Appendix E LLM Sources
OpenAI: GPT-4o, GPT-4-turbo, GPT-4o-mini
https://platform.openai.com/docs/api-reference/introduction
Anthropic: Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku
https://docs.anthropic.com/en/api/getting-started
Google: Gemini 1.5 Pro
https://ai.google.dev/gemini-api/docs
Meta: Llama-3.1-405B
https://www.llmapi.com
Mistral AI: Mistral Large
https://docs.mistral.ai/api/
DeepSeek: DeepSeek-V3
https://api-docs.deepseek.com/zh-cn/
Alibaba: Qwen-Max
https://bailian.console.aliyun.com/
ZhipuAI: GLM-4-Plus
https://www.bigmodel.cn/dev/api/normal-model/glm-4

### v2

@@ 8 Conclusion
In this study, we introduce the KGDS task, which aims to create observer-centric summaries by integrating shared background knowledge with discussions. We establish the first benchmark for KGDS and propose a novel hierarchical evaluation framework with fine-grained, interpretable metrics. Our evaluation of 12 leading LLMs reveals significant challenges in background summary retrieval, generation, and opinion summary integration, with even the most advanced models achieving less than 69% average performance across both paradigms.
@@
Our KGDS benchmark is developed with careful consideration of ethical implications. The news articles used as shared background knowledge are publicly accessible through Google News and have been carefully reviewed to ensure they do not contain sensitive or harmful information. Explicit consent was obtained from all expert participants involved in discussion construction and data annotation, with a clear explanation of the research purpose and data usage protocols. Our evaluation methodology prioritizes factual accuracy and opinion fidelity to minimize the risk of hallucination propagation in real-world applications. Potential misuse risks, such as generating misleading summaries through improper background-discussion combinations, are mitigated through technical safeguards in our released code and explicit usage guidelines. Researchers using our KGDS benchmark should adhere to responsible AI principles, especially when applying similar techniques to sensitive domains like healthcare or legal discussions. We provide the annotation and evaluation costs as:
@@
Annotation Cost for KGDS BenchMark.
Our annotation process is multi-step, fine-grained, and highly complex. For each sample in our benchmark, the average annotation time is hours (including news reading and understanding, discussion construction, supporting paragraphs, key supporting atomic facts, nonsupporting atomic facts, and clear atomic opinions annotation). We pay each annotation expert a wage of $ per hour (above the minimum wage standard), and the total annotation cost for the KGDS benchmark is $ ( hours per sample $ per hour experts per sample samples $). The evaluation cost of API calls is detailed in Table 6.
@@ Acknowledgement
This research was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081) and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18). We gratefully acknowledge the research funding provided by Fanyu AI Laboratory. We thank the anonymous reviewers for their constructive feedback that helped improve this work.
Appendix A Paradigm-Level Performance Visualization
Figure 6 and Figure 8 illustrate the overall performance of each model under the EBS-AOS and ABS-AOS paradigms, respectively. Figure 9 presents the average performance of each model across both paradigms. Figure 7 reveals the cross-paradigm stability of each model by quantifying the performance gap between the two paradigms.
Appendix B Further Analysis
LLMs are better retrievers than generators for background summary.
As presented in Table 3, all LLMs exhibit higher metrics compared to their counterparts. This gap between retrieval and generation indicates that LLMs possess stronger in-context recognition capabilities for coarse-grained paragraphs than fine-grained facts.
Cross-paradigm stability of LLMs.
By analyzing the performance gap between the two summarization paradigms (see Figures 6, 7, and 8 in Appendix A), we find that different LLMs excel in different paradigms. For example, Claude 3.5 Sonnet exhibits a significant gap ( ), while Claude 3 Opus demonstrates a relatively smaller gap ( ), indicating stronger stability.
LLMs perform better with EBS-AOS than ABS-AOS for KGDS.
As presented in Table 3, all LLMs achieve superior in the EBS-AOS paradigm compared to ABS-AOS, due to the more complete and accurate background summaries and higher-quality opinion summaries. Therefore, we suggest prioritizing EBS-AOS in real-world KGDS for more effective implementation.
Appendix C Detailed Formulas of Evaluation Metrics
C.1 ABS Evaluation Metrics
Let denote the LLM-generated ABS, and let and represent the sets of key supporting and nonsupporting atomic facts, respectively. We define the fact verification function as:
where if the fact can be inferred from , and otherwise.
KSAF Recall.
This metric quantifies the coverage of key supporting atomic facts in the ABS:
KSAF Precision.
This metric measures the focus of the ABS on key supporting atomic facts:
KSAF F1-score.
This metric evaluates the overall quality of the ABS by comprehensively considering both coverage and focus:
C.2 AOS Evaluation Metrics
Let denote the LLM-generated AOS, and let represent the set of clear atomic opinions. We define the opinion verification function as:
where if the opinion can be inferred from , and otherwise.
CAO Recall.
This metric quantifies the coverage of clear atomic opinions in the AOS:
Appendix D Detailed Benchmark Statistics
@@
Among the paragraphs from original articles, the expert annotation consistency rate is . A total of paragraphs are retained for the final news articles, while are identified as ambiguous paragraphs and removed. Of the retained, are annotated as supporting and as nonsupporting.
@@
Among the atomic fact units decomposed from the supporting paragraphs, are consistently annotated as key supporting atomic facts, as non-key atomic facts, and as intra-repetitive atomic facts.
Among the atomic fact units from the nonsupporting paragraphs, are automatically annotated as nonsupporting atomic facts, as masked conflicting facts, and as intra-repetitive atomic facts.
@@
A total of clear atomic opinions are expert manually annotated. Of these, contain clarified implicit references, while do not require clarification. Within the opinions, there are a total of written clarified implicit references.
@@ Appendix E LLM Sources
OpenAI111111https://platform.openai.com/docs/api-reference/introduction: GPT-4o, GPT-4-turbo, GPT-4o-mini
Anthropic121212https://docs.anthropic.com/en/api/getting-started: Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.5 Haiku
Google131313https://ai.google.dev/gemini-api/docs: Gemini 1.5 Pro
Meta141414https://www.llmapi.com: Llama-3.1-405B
Mistral AI151515https://docs.mistral.ai/api/: Mistral Large
DeepSeek161616https://api-docs.deepseek.com/zh-cn/: DeepSeek-V3
Alibaba171717https://bailian.console.aliyun.com/: Qwen-Max
ZhipuAI181818https://www.bigmodel.cn/dev/api/normal-model/glm-4: GLM-4-Plus
Appendix F Prompts and Instructions
F.1 Unified Parameter Settings
In this work, all 12 evaluated LLMs use consistent parameter settings for prompts and instructions: max_token=4096 and temperature=0. No other default parameters are modified.
F.2 Atomic Fact Decomposition Instruction
This instruction is used during the benchmark construction phase to guide a large language model in decomposing news paragraphs into fine-grained, atomic fact units. The instruction specifies three core principles for decomposition: indivisibility, independence, and declarativity. This ensures that the resulting fact units are minimal in granularity and informationally complete (see Figure 10).
F.3 Structured Prompts
These prompts serves as the core prompt for evaluating the performance of different LLMs on the KGDS task. It provides the models with standard inputs, including the shared background knowledge and the knowledge-grounded discussion, and clearly defines the objectives and requirements for the two sub-summaries under both summarization paradigms. See Figures 11 and 12 for EBS-AOS and ABS-AOS paradigms, respectively.
F.4 Self-Reflection Instructions
These instructions are used in the multi-turn self-reflection evaluation setting. After the model generates its initial summary, these instructions guide it to critically review its output, checking whether the generated background and opinion summaries strictly adhere to the task definitions. The instructions require the model to provide a detailed chain of thought in a step-by-step manner and, based on this reflection, generate a refined summary. This process is designed to explore and evaluate the model’s self-correction capabilities on the KGDS task. See Figures 13 and 14 for the EBS-AOS and ABS-AOS paradigms, respectively.
F.5 Verification Prompts
These prompts are a core component of our evaluation framework, used to automatically verify the quality of the generated summaries. We utilize a powerful verifier LLM (GPT-4o in our work) to execute these prompts, determining whether the model-generated summary contains the key information from the gold standard. Specifically, the fact verification prompt (see Figure 15) is used to assess the recall and precision of the abstractive background summary against key supporting atomic facts, while the opinion verification prompt (see Figure 16) evaluates the coverage of the abstractive opinion summary against clear atomic opinions.
F.6 Error Detection Instruction
This instruction is used for the fine-grained error analysis of the abstractive opinion summary. When the verifier determines that the summary has failed to cover a gold-standard clear atomic opinion, this instruction is invoked to attribute the failure to one of five predefined, specific error types (e.g., opinion misattribution, implicit reference incorrectly clarified). This approach allows us to gain deep insights into the specific weaknesses of each model in integrating opinions and clarifying references. We provide the instruction in Figure 17.